OpenStudy (anonymous):

Help me with story problem A regular pentagon has an apothem of 3.2 m and and area of 37.2 m. What is the length of one side of the pentagon?

OpenStudy (e.mccormick):

OK. How far did you get? What parts of this so you understand or not understand?

OpenStudy (anonymous):

That's a bad story. :c

jimthompson5910 (jim_thompson5910):

A = (ap)/2 37.2 = (3.2*p)/2 37.2*2 = 3.2p keep going to solve for p

jimthompson5910 (jim_thompson5910):

once you have the perimeter p, use it to find the side length s = p/n s = p/5

OpenStudy (anonymous):

The area of a regular polygon can be found by doing: \(\frac{1}{2}\text{apothem}*\text{perimeter}\) See if you can use that to find the perimeter.

OpenStudy (anonymous):

Got the answer guy THX